Electric Light tested a sample of 20 lightbulbs which produced a mean lifetime of 114 hours with a standard deviation of 4.7 hours. Construct 98% confidence interval for the true mean lifetime of the lightbulbs.
Answer #1

Answer:

Given,

Sample n = 20

Mean = 114

Standard deviation = 4.7

alpha = 1 - 0.98 = 0.02

alpha2 = 0.01

degree of freedom = n - 1

= 20 - 1

= 19

t(alpha/2,df) = t(0.01 , 19) = 2.54

Interval = xbar +/- t*s/sqrt(n)

substitute values

= 114 +/- 2.54*4.7/sqrt(20)

= 114 +/- 2.67

= (114 - 2.67 , 114 + 2.67)

= (111.33 , 116.67)

Interval = (111.33 , 116.67)